OpenStudy (summersnow8):

I just need help solving c. the answer is supposedly 731 N During the delivery phase of fastball pitch, the arm internally rotates at the shoulder. The angular velocity of this internal rotation peaks at 120 rad/s. At this instant, the elbow angle is 90 degrees, so the angular velocity of the forearm is also 120 rad/s. The baseball in the pitcher's hand is 35 cm from this axis of rotation through the shoulder joint. At this instant the linear velocity of the baseball is 45 m/s.

OpenStudy (anonymous):

sorry i thought i knew this but i dont sorry

OpenStudy (summersnow8):

a) How much of the baseball's total linear velocity is due to the 120 rad/s angular velocity of the forearm? --> 120 rad/s x .35 m = 42 m/s b) What is the centripetal acceleration of the baseball at this instant? \[(120 rad/s)^{2} \times (.35 m) = 5040 m/s^{2}\] c) How large is the force exerted by the pitcher on the baseball to cause this acceleration? The baseball's mass is 145 g.

OpenStudy (anonymous):

F=m * a F = 0.145 * 5040 = 730.8N